### v1

# Balancing Act: Distribution-Guided Debiasing in Diffusion Models

Abstract
Diffusion Models (DMs) have emerged as powerful generative models with unprecedented image generation capability. These models are widely used for data augmentation and creative applications. However, DMs reflect the biases present in the training datasets. This is especially concerning in the context of faces, where the DM prefers one demographic subgroup vs others (eg. female vs male). In this work, we present a method for debiasing DMs without relying on additional data or model retraining. Specifically, we propose Distribution Guidance, which enforces the generated images to follow the prescribed attribute distribution. To realize this, we build on the key insight that the latent features of denoising UNet hold rich demographic semantics, and the same can be leveraged to guide debiased generation. We train Attribute Distribution Predictor (ADP) - a small mlp that maps the latent features to the distribution of attributes. ADP is trained with pseudo labels generated from existing attribute classifiers. The proposed Distribution Guidance with ADP enables us to do fair generation. Our method reduces bias across single/multiple attributes and outperforms the baseline by a significant margin for unconditional and text-conditional diffusion models. Further, we present a downstream task of training a fair attribute classifier by rebalancing the training set with our generated data. Project Page
1 Introduction
Figure 1: a) Random sampling from a pretrained DM [33] generates images with gender imbalance. b) Proposed method takes a user-defined reference attribute $\mathbf{p^{a}_{ref}}$ and performs distribution guidance on a pretrained DM. c) Sampling with distribution guidance results in fair generation that follow user define $\mathbf{p^{a}_{ref}}$.
Recent advancements in Diffusion Models (DM) [11, 33, 42] have garnered much interest among researchers in evaluating the quality of the generated content. These models are not only used to generate realistic content but also to augment real datasets [45, 24] for downstream tasks. However, these popular models have been found to synthesize biased content with respect to multiple demographic factors like gender, race, age, etc., which can be detrimental to society once these models are deployed in the real world [29, 21, 34]. The problem is largely caused by the images used to train these models, as the outputs of these models are governed by these training datasets [29, 22]. Effects of such harmful biases have been shown by multiple recent works involving studies on not only DMs [29, 21, 34], but also GANs and other generative models [22, 5, 12]. In fact, Perera et al. [29] show that unconditional DMs–even when trained with balanced data–amplify racial biases, leading to the generation of more white-skinned faces than dark-skinned ones. Biases in the generated data are even more evident in large text-to-image DMs, e.g., models mostly tend to generate a specific gender with a given profession (like male and doctor) [23, 52, 39].
Existing works on debiasing, either require a reference dataset  [5, 52] and/or allow retraining of the model [5, 50, 51]. On the contrary, our work aims to mitigate biases in both unconditional and conditional DMs, enabling fair generation without model retraining. We propose a setting where we provide a reference attribute distribution ($\mathbf{p^{a}_{ref}}$) for sensitive attribute $a$ and query the DM to sample images following $\mathbf{p^{a}_{ref}}$. This framework allows for adapting any existing DM to a pre-defined distribution. E.g., a user can provide $\mathbf{p^{a}_{ref}}$ as a uniform histogram of sensitive attributes to generate balanced attribute distribution. We believe defining $\mathbf{p^{a}_{ref}}$ provides just enough information to condition the DM for fair generation. This is an extremely practical setting for debiasing and is particularly important for large text-to-image DMs [33, 36, 32] where retraining or fine-tuning is computationally intensive.
One plausible approach for fair generation is to guide every generated sample with attribute classifiers following classifier guidance [8]. However, such a framework, though simple, is overly restrictive as it requires presetting and enforcing attributes for each sample individually (which we call sample guidance). Such constraints during denoising result in inferior generation quality, as discussed in Sec.4.1. Instead, we propose to jointly denoise a whole batch of samples and guide the process with $\mathbf{p^{a}_{ref}}$ (which we call distribution guidance). Specifically, we push the generated batch attribute distribution $\mathbf{{p^{a}_{\theta}}}$ and $\mathbf{p^{a}_{ref}}$ close to each other during the reverse process. Distribution guidance provides more flexibility to each sample during generation as it does not enforce a preset of attributes on a sample basis. This results in fair generation without sacrificing the generation quality.
A major challenge for guidance-based conditioning is that it requires separate image classifiers for each noise scale of the diffusion process. To overcome this, we propose to perform guidance in semantically rich features space - h-space [18] of DMs. Specifically, we train an Attribute Distribution Predictor (ADP) that predicts attribute distribution directly from the h-space features. As ADP is trained on rich and discriminative h-space features, it is - a) implemented as a linear layer, b) requires minimal training data, and c) fast training and inference. Finally, during inference, we steer the h-space representation by matching the predictions from ADP to $\mathbf{p^{a}_{ref}}$.
We extensively evaluate our proposed method for the fair generation of single and multi-attribute cases for face generation. Additionally, we present the results of our method on Stable Diffusion [33], a large text-to-image DM. Further, as downstream application train debiased attribute classifiers by augmenting the training data for minority subgroups.
The major contributions of this work are the following:
  1. 1.
    A novel setting for debiasing existing DMs without retraining given a reference attribute distribution.
  2. 2.
    Distribution guidance to condition the reverse diffusion process on a reference attribute distribution.
  3. 3.
    Propose guidance in the intermediate features of diffusion network (h-space), which leads to data-efficient training and fast generation.
2 Related Works
Biases in Generative Models. While generative models like Generative Adversarial Networks and Diffusion Models have become the de-facto tools for image generation in recent times, studies show that they are not free of biases [29, 21, 34, 22, 13]. Perera et al.  [29] show that unconditional diffusion models amplify the biases in the training data with respect to gender, race, and age. Luccioni et al.  [21] identify and quantify social biases in images generated by popular text-to-image models like DALL-E 2, and Stable Diffusion v1.4 and 2. Maluleke et al.  [22] study racial biases in GANs and find that GANs mimic the racial distribution in the training data.
Debiasing generative models by retraining. This line of work focuses on mitigating biases in generative models by retraining them [27, 50, 47, 38, 51, 5, 44, 46]. Some of these works [50, 38, 47] assume knowledge of the labels of the sensitive attribute and then debias the models such that there is no correlation between the decision attribute and the sensitive attribute. IMLE-GAN [51] ensures coverage of minority groups by combining GAN adversarial training with Implicit Maximum Likelihood Estimation (IMLE) [20]. Another body of works employs a balanced unlabelled reference dataset to ensure unbiased generations [5, 44, 46]. Choi et al. [5] use a density-ratio based technique to identify the bias in datasets via the reference dataset, and learn a fair model based on importance reweighting with the help of both the original biased and the reference dataset. To capture the distance between the small reference data and the generated data, Um et al. [46] use the LeCam Divergence [19]. On the other hand, Teo et al. [44] introduce a transfer learning approach to solve this problem by training the model on the biased dataset first and then adapting the model to the reference set.
Debiasing generative models without training. As training of GANs and DMs can be resource-consuming, many methods prefer fair generation of images without explicit training [31, 43, 24, 12]. MaGNET [12] aims to produce uniform sampling on the learned manifold of any generative model like GANs or VAEs, while Ramaswamy et al.  [31] and Tan et al.  [43] manipulate the latent space of GANs to generate balanced outputs. GANDiffFace [24], on the other hand, generates balanced synthetic for face recognition by first generating high-quality images from different demographics using GANs and then finetuning Stable Diffusion [33] using DreamBooth [35] to generate more images of such identities with different poses, expressions, etc. Multiple works attempt to mitigate biases in vision-language models and text-conditioned diffusion models as well [6, 52, 41, 53, 3]. Chuang et al.  [6] debias the text embedding using a calibrated projection matrix and shows unbiased generations without any additional training or data. However, debiasing the unconditional DMs has received less attention, which is the main focus of this work.
Guidance in Diffusion Models. One of the primary techniques to condition the diffusion model is to guide the reverse diffusion model with the gradients of additional network  [8]. GLIDE [28] used CLIP [30] based guidance for open world caption to image generation. Guidance is used for image-to-image translation [49], counterfactual generation [37]. However, guidance in its original form requires retraining of guiding networks on the noisy data from scratch. Few works overcome this by learning a mapping function from the diffusion feature space to sketches for sketch-guidance  [48] and a universal guidance that repurposes pretrained networks for guidance  [1].
3 Method
We assume a setting where we are given a pretrained DM trained on biased data and a reference distribution of the sensitive attributes $\mathbf{p^{a}_{ref}}$. Our goal is to generate data from the DM, whose generated attribute distribution $\mathbf{p^{a}_{\theta}}$ best approximates the reference $\mathbf{p^{a}_{ref}}$ without retraining.
The key idea is to jointly guide the denoising of a batch of samples such that $\mathbf{p^{a}_{ref}}$approx $\mathbf{p^{a}_{\theta}}$. Directly computing $\mathbf{p^{a}_{\theta}}$ in a closed form is intractable. Instead, we train an Attribute Distribution Predictor, a small network that maps the intermediate batch features of a diffusion network to an estimate of attribute distribution $\mathbf{\hat{p}^{a}_{\theta}}$.
Figure 2: Sample guidance vs Distribution guidance. a) After a few steps of denoising ($t=\tau$), the generated samples have learned some discriminative features for gender. Sample guidance enforces a preset attribute state (male/female) to all the samples in the batch. This is stringent and can result in enforcing samples with dominant female features to be changed. However, in distribution guidance, only those samples are converted, which are easy to convert and are close to the decision boundary. b) Visualization of the generated samples with different guidance strengths. It can be observed that sample guidance is not able to change the attribute in some cases, and increasing guidance strength results in collapse.
3.1 Preliminary
Diffusion models have emerged as a powerful family of generative models trained to learn the data distribution by gradual denoising from a Gaussian distribution. Starting from a clean point $\mathbf{x_{0}}$, and a set of scalar values $\{\alpha_{t}\}_{t=1}^{T}$, applying $\mathbf{t}$ steps of the forward diffusion process yields a noisy data point, where $\bar{\alpha_{t}}$ = $\prod_{i=1}^{t}\alpha_{i}$
$\mathbf{x_{t}}=\sqrt{\bar{\alpha_{t}}}\mathbf{x_{0}}+(\sqrt{1-\bar{\alpha_{t}}})\epsilon,\epsilon\approx\mathcal{N}(0,I)$ (1)
A diffusion model is learned as a neural network $\epsilon_{\theta}$ that predicts the noise from given $\mathbf{x_{t}}$ and $\mathbf{t}$. The reverse process takes the form $q(\mathbf{x_{t-1}}|\mathbf{x_{t}},\mathbf{x_{0}})$, which is parameterized as a Gaussian distribution. In this work, we consider DDIM [42] sampling which first computes a clean data point $\mathbf{\hat{x}_{0}}$ and then sample $\mathbf{x_{t-1}}$ from $q(\mathbf{x_{t-1}}|\mathbf{x_{t}},\mathbf{\hat{x}_{0}})$.
Classifier guidance is proposed to condition a diffusion model on class labels with the help of a pretrained classifier [8]. Specifically, a classifier $f_{\phi}(c|\mathbf{x_{t}},t)$ is trained on noisy images at different noise scales, where $c$ is the class label. The gradients of the classifier are then used to guide the diffusion sampling process towards the class label $c$. Concretely, the classifier guidance performs sampling by updating the noise prediction $\epsilon_{\theta}(\mathbf{x_{t}},\mathbf{t})$.
$\hat{\epsilon}_{\theta}(\mathbf{x_{t}},t)=\epsilon_{\theta}(\mathbf{x_{t}},t)-\sqrt{1-\alpha_{t}}\nabla_{\mathbf{x_{t}}}\log f_{\phi}(c|\mathbf{x_{t}},t)$ (2)
3.2 Classifier guidance for debiasing
A promising approach is to leverage pretrained attribute classifiers to guide toward balanced generation. Assuming a reference attribute distribution $\mathbf{p^{a}_{ref}}$ for a binary attribute $\mathbf{a}$ (e.g., gender) and corresponding attribute classifier ${f}_{\phi}$. We parameterize $\mathbf{p^{a}_{ref}}$ as a Bernoulli distribution with parameter $\mathbf{r}$ denoting fraction of males samples. To generate samples following $\mathbf{p^{a}_{ref}}$, we can randomly select $\mathbf{Nr}$ samples from the batch size of $\mathbf{N}$ and guide them towards male class using predictions from ${f}_{\phi}$. Similarly, for the remaining $\mathbf{N(1-r)}$ samples, we guide them towards female class. In practice, such a sample guidance follows $\mathbf{p^{a}_{ref}}$ up to some extent, but results in inferior sample quality (see Fig. 2b).
Insight. 1: Attempting to steer the generation of individual samples towards a pre-defined attribute class is overly restrictive and leads to inferior generation quality.
Remark: Sample guidance requires enforcing a preset attribute state (male / female) to each sample of a batch during the denoising process, which is too stringent and results in distorting the outputs. Given an intermediate time-stamp $\tau$, the samples in the earlier stages ($t>\tau$) of reverse diffusion process are close to the noise space, resulting in the poor classifier of the corresponding stages. The guidance becomes effective only in the later stage of denoising ($t<\tau$). However, till timestep $\tau$, some facial features are already formed for the samples [25]. Hence, enforcing a preset attribute state at later stages ($t<\tau$) is restrictive and results in sample collapse (Fig. 2-b)). For e.g., for gender attribute, if a sample has formed dominant female features till $\tau$, enforcing it to be $male$ is overly restrictive, and classifier guidance distorts the generated sample.
3.3 Distribution Guidance
We propose an alternate guidance strategy termed Distribution Guidance for fair generation, which provides more flexibility to modify the attribute states during the denoising process. The key idea is to jointly denoise a batch of samples $\mathbf{x_{T}^{[1:N]}}$, with the guidance from reference attribute distribution $\mathbf{p^{a}_{ref}}$. To realize distribution guidance, we define a differentiable distribution prediction function $g_{\psi}$ that maps the batch samples to an estimate of generated attribute distribution $\mathbf{\hat{p}^{a}_{\theta}}$. We learn the function $g_{\psi}$ over an intermediate feature space - h-space [18] of denoising network instead of image space for efficiency (Sec.3.4). Hence, the batch estimate is given by, $\mathbf{\hat{p}^{a}_{\theta}}=g_{\psi}(\mathbf{h_{t}^{[1:N]}},\mathbf{t})$, where $\mathbf{h_{t}^{[1:N]}}$ is the bottleneck U-Net feature representation of the batch samples $\mathbf{x_{t}^{[1:N]}}$. Further, we define a loss function $\mathcal{L}(.)$ that measures the similarity of two distributions. During denoising we guide the batch of samples to bring $\mathbf{\hat{p}^{a}_{\theta}}$ closer to the reference, i.e., $\mathcal{L}(\mathbf{\hat{p}^{a}_{\theta}},\mathbf{p^{A}_{ref}})\approx 0$. This can be easily integrated into the reverse diffusion process as an extension of classifier guidance by modifying Eq.2.
Insight.2: Distribution guidance provides flexibility to batch samples to easily modify the attributes to match the defined attribute distribution.
Remark: As distribution guidance does not require presetting for attribute states for each sample, it gives more flexibility during the generation as long as $\mathbf{\hat{p}^{a}_{\theta}}$ follows $\mathbf{p^{a}_{ref}}$. E.g., in our running example of gender attribute, only those samples will change gender after $t=\tau$, which are close to the decision boundary. In contrast to sample guidance, where random samples are forced to change the attribute state (Fig.  2) resulting in inferior quality, distribution guidance is able to generate fair samples with good quality. Observe that distribution guidance covers the samples that are close to the decision boundary for gender classification, resulting in good outputs.
A major design decision for implementing distribution guidance is the selection of function $g_{\psi}$. The conventional approach is to train multiple attribute classifiers at each level of noise in the image space. Instead, we propose to leverage intermediate semantic features from the diffusion model itself and repurpose them for the task of attribute classification. This framework is super efficient as compared to image space classifiers.
3.4 Guidance in the H-space
Diffusion models, although trained for image generation, learn semantically rich representations in the internal features of denoising network $\mathbf{{\epsilon_{\theta}}}$. These representations have been successfully used for segmentation [2] and classification [26]. Motivated by this, we ask can we use these internal features for training guidance classifiers? To answer this, we take features from the bottleneck layer of the diffusion U-Net, termed as the h-space [18] and attach a linear head for classification. The trained classifiers achieve good classification performance across multiple attributes, as shown in Fig. 3. Hence, we use the trained $h-space$ classifiers to realize $g_{\psi}$. Specifically, we train a network Attribute Distribution Predictor that maps the batch h-space features to an estimate of attribute distribution $\mathbf{\hat{p}^{a}_{\theta}}$.
Figure 3: Classification accuracy for linear h-space classifiers and ResNet-18 image space classifiers trained on $2K$ training examples. h-space classifiers are data efficient and achieve superior performance even with a linear layer
Figure 4: Distribution guidance in the h-space. For a given batch $\mathbf{x_{t}^{[1:N]}}$, we extract the intermediate h-space representation $\mathbf{h_{t}^{[1:N]}}$ and pass it through ADP to obtain attribute distribution $\mathbf{\hat{p^{a}_{\theta}}}$. Guidance updates $\mathbf{h_{t}^{[1:N]}}$ by backpropagating the derivative of loss.
Attribute Distribution Predictor (ADP) is realized via a linear attribute classifier conditioned on the diffusion time step $\mathbf{t}$. Given a batch of generating samples $\mathbf{x_{t}^{[1:N]}}$, we extract the corresponding h-space features $\mathbf{h_{t}^{[1:N]}}$ = $\epsilon^{E}_{\theta}(\mathbf{x_{t}^{[1:N]}},\mathbf{t})$, where $\epsilon^{E}_{\theta}$ is the encoder of U-Net network. Next, we pass the batch $\mathbf{h_{t}^{[1:N]}}$ to the attribute classifier and obtain a batch of softmax predictions. The softmax predictions are aggregated per class to obtain the attribute distribution estimate $\mathbf{\hat{p}^{a}_{\theta}}$. Finally, we update the intermediate h-vectors with the gradients of the distribution matching loss $\mathcal{L}$, where $\mathbf{\gamma}$ is the guidance strength parameter:
$\mathbf{\tilde{h}_{t}^{[1:N]}}=\mathbf{h_{t}^{[1:N]}}$ $-\gamma*\nabla_{\mathbf{h_{t}^{[1:N]}}}\mathcal{L}(\mathbf{\hat{p}^{a}_{\theta}},\mathbf{p^{a}_{ref}})$
$\mathbf{\hat{p}^{a}_{\theta}}=$ $g_{\psi}(\mathbf{h_{t}^{[1:N]}},\mathbf{t})$
Finally, we obtain the distribution guided noise predictions $\tilde{\epsilon}$ by passing the $\mathbf{h^{[1:N]}_{t}}$ through the U-Net decoder $\epsilon^{D}_{\theta}$, i.e., $\tilde{\epsilon}(\mathbf{x^{[1:N]}_{t}},\mathbf{t})=\epsilon^{D}_{\theta}(\mathbf{h^{[1:N]}_{t}},\mathbf{t})$. The predicted noise is then used to update batch $x^{[1:N]}_{t-1}$ using DDIM [42].
Insight.3: H-space guidance is extremely effective and efficient as compared to image space guidance.
Remark. As compared to conventional image space guidance, guidance in h-space has multiple advantages: i) it requires only a set of linear layers to implement the classifiers, ii) it is fast to backpropagate during guidance as compared to image models, iii) highly data-efficient and can be trained with only a few thousand examples due to semantically rich h-space. In the experiments section, we compare these properties of the h-space guidance.
4 Experiments
In this section, we first discuss the evaluation metrics, followed by the implementation details. We take face datasets as the subject of study, as they are subject to very high demographic biases. We compare our method with other inference time debiasing approaches for both single-attribute and multi-attribute debiasing. Further, we present detailed ablations to validate design choices. Finally, we show the generalization of our debiasing approach on large text-to-image generation model Stable Diffusion [33].
4.1 Evaluation Metrics
A fair generative model is evaluated on two grounds: image quality and fairness. Therefore, we begin by discussing the metrics we use to measure these two attributes of the generated images.
Fairness. The primary goal of this paper is to generate images based on the reference distribution $\mathbf{p_{ref}^{a}}$. We follow Choi et al.  [5] to define the Fairness Discrepancy (FD) metric. Given an attribute $\mathbf{a}$, we assume access to a high-accuracy classifier for $\mathbf{a}$ (denoted as $\mathcal{C}_{a}$), and using the predictions from the latter, we compute the following [5, 44]:
$|\bar{p}-\mathbb{E}_{\mathbf{x}\sim p_{\theta}(\mathbf{x})}(\mathbf{y})|_{2}$
where $\mathbf{y}$ is the softmax output of the classifier $\mathcal{C}_{a}(\mathbf{x})$, $\bar{p}$ is a uniform vector of the same dimension as $\mathbf{y}$, $p_{\theta}$ is the distribution of the generated images. The lower the FD score, the closer the distribution of the attribute values is to the uniform distribution – i.e., the generated images are fairer with respect to attribute $\mathbf{a}$.
Image Quality. To measure the quality of generation, we follow the standard Fréchet Inception Distance (FID) [10]. We compute this metric with the help of an attribute-wise balanced dataset sampled from the original training data.
4.2 Implementation Details
Training h-space classifiers. We start with creating a paired dataset $\mathcal{D}^{h}_{clf}$ of h-vector and attribute labels. Specifically, we take a subset $\mathcal{D}$ of the CelebA-HQ [15] dataset and obtain attribute labels for the images using the pretrained attribute classifier $\mathcal{C}_{a}$. Next, we embed image $\mathbf{I^{i}}\in\mathbf{D}$ to obtain the corresponding h-space representation $\mathbf{\mathcal{H}^{i}}=\{\mathbf{h^{i}_{t}}\}^{t=T}_{t=0}$ using DDIM [42] inversion. This yields a labelled dataset $\mathcal{D}^{h}_{clf}$ with pairs $(\mathcal{H}^{i},\mathbf{y^{i}})$, where $\mathbf{y^{i}}=\mathcal{C}_{a}(I^{i})$ is the predicted attribute label for image $\mathbf{I^{i}}$ (e.g. male / female). Next, we train h-space attribute classifiers $\mathcal{C}^{h}_{a}(\mathbf{h_{t}},t)$ as a linear head over $\mathbf{h_{t}}$ and conditioned on time $t$ (ignored sample index $i$ for brevity). These obtained classifiers generate high-accuracy attribute predictions, as shown in Fig. 3. Further details about the dataset and classifiers are provided in a supplementary document.
Figure 5: Balancing generated data across gender and eyeglass attributes with various guidance strategies. Minority group $male$ and $eyeglasses$ are marked. Image-space Universal Guidance and Sample Guidance results generate imbalanced and poor-quality images. For gender, Sample Guidance is able to generate only a few males while maintaining quality. The proposed Distribution Guidance balances (close to the ratio of $0.50:0.50$) both the attributes and generates high-quality images.
Distribution guidance. To perform distribution guidance, we realize the distribution prediction function, ADP with h-space classifiers $\mathcal{C}^{h}_{a}$. Specifically for a generating batch $\mathbf{x_{t}^{[1:N]}}$, we obtain the corresponding $h-space$ representation $\mathbf{h_{t}^{[1:N]}}$ and obtain a set of attribute predictions $\mathbf{\hat{y}^{[1:N]}_{t}}$ from classifiers $\mathcal{C}^{h}_{a}$. Finally, we add all the softmax values for all the $N$ images in the batch for each class to obtain the estimate of $\mathbf{p^{a}_{\theta}}$. We use Chi-square distance as the loss $\mathcal{L}$ with the reference distribution $\mathbf{p^{a}_{ref}}$.
Diffusion Model Architecture. We evaluated our proposed debiasing method for fair generation from two state-of-the-art pretrained DMs: a) an unconditional DM, P2  [4], trained on CelebA-HQ dataset, and b) a text conditional DM StableDiffusion v1.5  [33] trained on the LAION [40] dataset. Both these models have exceptional image generation quality; however, they have significant bias concerning the sensitive face attributes, as shown in the following subsections.
4.3 Evaluation and Comparisons.
We compare our proposed method against two techniques for guidance-based generations for DMs and a state-of-the-art sampling-based technique for debiasing generative models, MagNet [12].
Sample guidance. We use classifier guidance as explained in Sec. 3.2 in the h-space using the trained h-space attribute classifiers $\mathcal{C}_{a}^{h}$. Such guidance requires presetting an attribute state for each generating image and pushing the trajectory to change the attribute state.
Universal guidance [1] performs guidance in the image space but uses pretrained classifiers trained on the clean image. This resolves the additional requirement of training image classifiers on noisy images. The key idea is to use the DDIM scheduler and predict the approximation of $\mathbf{\hat{x_{0}}}$ from noise image $\mathbf{x_{t}}$, and pass it through pretrained image classifier $\mathcal{C}_{A}$. However, this process has two shortcomings. Firstly, it is slow as we have to backpropagate the gradients from the image classifier. Experimentally, we found the h-space guidance to be $7$ times faster than the image guidance. Secondly, it performs poorly at the early stage due to an inaccurate approximation of $\mathbf{x_{0}}$. We used the same training set of $|\mathcal{D}_{clf}|=2K$ to train the h-space and image space attribute classifiers for a fair comparison.
Table 1: Evaluation of balanced generation for single attribute
Gender Race Eyeglasses Metric FD FID FD FID FD FID Random Sampling 0.178 54.59 0.334 60.01 0.251 75.21 Universal Guidance  [1] 0.193 52.10 0.377 93.42 0.189 64.55 H-Sample Guidance 0.113 51.46 0.184 56.53 0.118 57.63 H-Distribution Guidance 0.049 50.27 0.113 52.38 0.014 51.78 Random sampling* 0.307 112.28 0.463 123.97 0.276 117.83 Magnet* 0.267 91.15 0.454 97.05 0.281 106.55
Quantitative evaluation. We evaluate our debiasing method for the single attribute case by generating balanced distributions of individual sensitive attributes - gender, eyeglasses, and race, in Tab. 1. For the scope of this paper, we consider all these attributes as binary, and hence expect a $1:1$ ratio of the sensitive attributes in the synthesized images (for e.g. $0.50$ fraction of males and $0.50$ fraction of females in case of gender). Specifically, we generate $10K$ images from each method per attribute and compute the metrics defined in Sec. 4.1. The proposed guidance method outperforms all the methods in terms of visual quality as measured by FID and bias generation measured by FD. Sample guidance achieves comparable FID for gender; however, higher FD indicates inferior debiasing. The tradeoff between FD and FID is discussed in detail in subsection 4.5. This supports our thesis that the distribution guidance provides enough flexibility during generation, resulting in high-quality outputs even with a high guidance scale.
Multiple attributes. Additionally, we apply our method for debiasing multiple attributes simultaneously in Tab. 2. Specifically, given two reference distributions $\mathbf{p^{a_{1}}_{ref}}$ and $\mathbf{p^{a_{2}}_{ref}}$, we add guidance from two pretrained attribute distribution predictors $g_{\psi_{1}}$ and $g_{\psi_{2}}$. For this experiment, we define both the reference distribution as uniform for each attribute ($50\%$-$50\%$ splits). The generated results follow the reference distribution, resulting in a balanced generation across attributes.
Table 2: Evaluation of balanced generation for multiple attribute
Gender + Race Eyeglasses + Race Gender + Eyeglasses Metric FD $\downarrow$ FID $\downarrow$ FD $\downarrow$ FID $\downarrow$ FD $\downarrow$ FID $\downarrow$ Random Sampling 0.256 60.68 0.292 89.14 0.214 70.97 Universal Guidance 0.283 71.84 0.264 91.54 0.157 80.57 H-Sample Guidance 0.241 59.78 0.135 67.87 0.079 52.03 H-Distribution Guidance 0.075 49.91 0.101 57.46 0.057 47.45
Comparison with MagNet [12]: MagNet is an unsupervised method enabling fair sampling from a pretrained model. They propose a method for uniform sampling on the image manifold to generate under-represented groups equally. We generated results by MagNet sampling from a StyleGAN2 [17] model trained on the FFHQ [16] dataset from their official codebase. We compare our method with MagNet [12] for the task of fair generation in Tab. 1. We used MagNet to generate $10K$ samples and compute the FD score. As the base model for MagNet is StyleGAN2, we cannot compare FID directly with our DM debiased results. However, we can observe that MagNet achieves similar FD scores to its baseline model and is ineffective in debiasing the three sensitive attributes: gender, race, and eyeglasses. However, our distribution guidance significantly improves the FD metric from the baseline DM model.
Qualitative evaluation. We present qualitative results for attribute balancing with respect to gender and eyeglasses in Fig. 5. We randomly sample $20$ starting noise and use individual guidance methods for debiasing. Without guidance, the DM mostly generates female faces for the gender attribute. Sample guidance and universal guidance, although increasing the number of males, collapses some images. On the other hand, Distribution guidance generates an almost equal number of males and females without affecting the generation quality. Moreover, the DM hardly generates images with eyeglasses for sample and universal guidance and also without any guidance. Our method, on the contrary, leads to highly balanced generations.
4.4 Generating Imbalanced Distributions
We tested our distribution guidance for generating imbalanced attribute distribution by providing skewed $\mathbf{p^{a}_{ref}}$ i) $0.20$ female and $0.80$ male ii) $0.10$ white race and $0.90$ black race. These two settings are extremely challenging given that male and black race are minority groups in the dataset. We present qualitative results for both the settings in Fig.6, where our distribution guidance is able to generate the defined distribution with majority males. Note, we have binarized the race attribute as black and whites for simplicity, and hence in the generation brown race is considered under black category. We report quantitative metrics in Tab. 3, where the proposed method can achieve good FD scores with the same FID.
Figure 6: Generating imbalanced data with two non-uniform reference distribution - i) 0.9B-0.1W, ii) 0.8M-0.2F. Minority classes are marked. Distribution guidance generates high-quality images and closely follows the skewed reference distribution.
Table 3: Distribution guidance for imbalanced generation
0.20F - 0.80M 0.10W - 0.90B Method FD FID FD FID Random Sampling 0.478 72.26 0.734 77.63 H-Distribution Sampling 0.168 51.65 0.325 53.80
4.5 Ablations
Guidance Strength $\gamma$. We ablate over guidance strength parameter $\gamma$ for both the distribution guidance and sample guidance in Fig. 7. Increasing $\gamma$ reduces the bias (better FD) at the cost of inferior image quality (increased FID). Guidance strength $1500$ achieved a good tradeoff between image quality and balancing attributes. Further, as compared to sample guidance, the distribution guidance achieves a better tradeoff for all the guidance strengths.
Figure 7: Ablation over guidance strength $\gamma$
Batch size. As we approximate $\mathbf{p^{a}_{\theta}}$ with a an estimate $\mathbf{\hat{p}^{a}_{\theta}}$ over a batch of size $N$, we ablate over different values of $N$ in Tab. 4. Intuitively, using a larger batch size yields a better estimate of $\mathbf{p^{a}_{\theta}}$. We have found that the $N=100$ works best for all our experiments.
Table 4: Ablation over batch size for gender balancing.
Batch size 10 25 50 75 100 FD $\downarrow$ 0.062 0.059 0.049 0.059 0.046 FID $\downarrow$ 50.54 51.51 50.86 50.98 51.64
Figure 8: Debiasing results on stable diffusion for gender attribute. The proposed distribution guidance can balance the gender attributes in the neutral prompt and generalizes well to prompts involving other professions e.g. firefighter, doctor.
4.6 Experiments on Stable Diffusion
We implemented our distribution guidance technique for debiasing a text-to-image generation model, Stable Diffusion (SD) v1.5 [33] concerning the gender attribute in this subsection. First, we generate a dataset $\mathcal{D}$, of $10K$ images from SD, with prompts ‘a photo of a male’ and ‘a photo of a female’ to generate a labeled dataset for training h-space classifiers. If we use the same CelebA-HQ dataset to train the h-space classifier, the guidance is ineffective due to a significant domain shift from the SD generations. Next, we obtain the corresponding labeled dataset $\mathcal{D}^{h}_{clf}$ in the h-space and train a gender classifier in the h-space. We used the trained classifier for distribution guidance following Sec. 3.4 to generate images with balanced gender.
It is observed that SD increases the gender bias when queried to generate certain professions (e.g. male and doctor) [52]. To this end, we implement our distribution guidance along the prompts ‘a photo of a doctor’ and ‘a photo of a firefighter’ to evaluate the effectiveness in this challenging setting. The qualitative results are shown in Fig.8, and the quantitative results are reported in Tab. 5. We can observe, as compared to the baseline (random generation without guidance), our method generates fairer gender distribution with both neutral prompts and prompts mentioning professions (e.g. ‘a photo of a firefighter’). Additional results on text-to-image models are present in supp. material.
Table 5: Balancing gender on Stable Diffusion [33] model.
Method Gender Doctor Firefighter FD $\downarrow$ FID $\downarrow$ FD $\downarrow$ FID $\downarrow$ FD $\downarrow$ FID $\downarrow$ Random Sampling 0.317 72.37 0.355 70.11 0.235 71.86 H-Sample Guidance 0.026 70.96 0.021 68.43 0.097 70.42 H-Distribution Guidance 0.024 70.69 0.015 67.36 0.093 69.41
4.7 Class-imbalance in attribute classification
We explore an important downstream application of our proposed approach in balancing minority classes by augmenting the under-represented classes with generated data. Specifically, we train a race classifier (labels obtained from $\mathcal{C}_{\mathcal{A}}$) on the CelebA-HQ [14] dataset. The race classifier is an ImageNet [7]-pretrained ResNet-18 [9] encoder, followed by 2 MLP layers and a classifier. We manually oversample the Whites and undersample the Blacks in the training dataset such that the imbalanced dataset consists of $10k$ samples of white people and $1k$ black people (we keep the genders balanced within a race class). Consequently, the model performs poorly on the minority class (i.e., Black) due to under-representation. Next, we augment (class-balanced) the training data by generating samples whose distribution is inversely proportional to the class counts in the training set to increase images of minority classes using our distribution guidance approach. This adds 9k images of only Black race, and the classifier trained on this balanced data performs significantly better (Tab. 6). Even when gender is balanced in both the classes, we observe a significant disparity in the accuracies for Male and Female Black samples in the vanilla classifier. However, our proposed method helps reduce the accuracy gap between Black males and Black females. This shows a potential application in generating ’class’ balanced datasets to train models for other downstream tasks, which can also mitigate bias.
Table 6: Group-wise accuracies of identical classifiers trained on existing and generated(balanced) data
Black female White female Black male White male Vanilla classifier 75.76 98.96 70.29 99.93 Balanced classifier 91.04 97.33 90.72 97.68
5 Discussion.
Limitations. Although our method performs guidance in the h-space, which is efficient compared to the image space guidance, it still requires additional training of h-space classifiers. Another limitation is reliance on accurate attribute classifiers to obtain labels for training h-space classifiers.
Future works. An important future work is extending distribution guidance beyond de-biasing for controlled generation and data augmentation. In the context of debiasing DMs, extending the proposed approach without needing an attribute classifier or labeled data.
6 Conclusion
In this work, we aim to mitigate biases from pretrained diffusion models without retraining - given only a desired reference attribute distribution. We propose a novel approach leveraging distribution guidance that jointly guides a batch of images to follow the reference attribute distribution. The proposed method is effective and results in both high-quality and fair generations across multiple attributes and outperforms sample guidance strategies based on conditioning each sample individually. Extensive experiments demonstrate the effectiveness of our method in balancing both single and multiple attributes on unconditional DMs and conditional text-to-image diffusion models. We believe such a setting of debiasing without retraining is practical, especially in today’s era of large-scale generative models.

### v2

@@
 female vs male).
-In this work, we present a method for debiasing DMs without relying on additional data or model retraining.
+In this work, we present a method for debiasing DMs without relying on additional reference data or model retraining.
 Specifically, we propose Distribution Guidance, which enforces the generated images to follow the prescribed attribute distribution.
@@
 Our method reduces bias across single/multiple attributes and outperforms the baseline by a significant margin for unconditional and text-conditional diffusion models.
-Further, we present a downstream task of training a fair attribute classifier by rebalancing the training set with our generated data.
-Project Page
+Further, we present a downstream task of training a fair attribute classifier by augmenting the training set with our generated data.
+Code is available at - project page.
 1 Introduction
@@
 These models are not only used to generate realistic content but also to augment real datasets [54, 29] for downstream tasks.
-However, these popular models have been found to synthesize biased content with respect to multiple demographic factors like gender, race, age, etc., which can be detrimental to society once these models are deployed in the real world [29, 21, 34].
+However, existing DMs have been found to synthesize biased content with respect to multiple demographic factors like gender, race, etc., which can be detrimental to society once these models are deployed in the real world [35, 26, 41].
 The problem is largely caused by the images used to train these models, as the outputs of these models are governed by these training datasets [35, 27].
-Effects of such harmful biases have been shown by multiple recent works involving studies on not only DMs [29, 21, 34], but also GANs and other generative models [22, 5, 12].
+Effects of such harmful biases have been shown by multiple recent works involving studies on DMs [35, 26, 41], GANs and other generative models [27, 6, 14].
 In fact, Perera et al.
@@
 On the contrary, our work aims to mitigate biases in both unconditional and conditional DMs, enabling fair generation without model retraining.
-We propose a setting where we provide a reference attribute distribution ($\mathbf{p^{a}_{ref}}$) for sensitive attribute $a$ and query the DM to sample images following $\mathbf{p^{a}_{ref}}$.
+We propose a practical setting where we provide a reference attribute distribution ($\mathbf{p^{a}_{ref}}$) for sensitive attribute $a$ and query the DM to sample images following $\mathbf{p^{a}_{ref}}$.
 This framework allows for adapting any existing DM to a pre-defined distribution.
-E.g., a user can provide $\mathbf{p^{a}_{ref}}$ as a uniform histogram of sensitive attributes to generate balanced attribute distribution.
+E.g., a user can provide $\mathbf{p^{a}_{ref}}$ as a uniform distribution for a given sensitive attribute to generate balanced attribute distribution.
 We believe defining $\mathbf{p^{a}_{ref}}$ provides just enough information to condition the DM for fair generation.
@@
 Distribution guidance provides more flexibility to each sample during generation as it does not enforce a preset of attributes on a sample basis.
+Intuitively, distribution guidance prioritizes transforming easier samples close to the decision boundary.
 This results in fair generation without sacrificing the generation quality.
 A major challenge for guidance-based conditioning is that it requires separate image classifiers for each noise scale of the diffusion process.
-To overcome this, we propose to perform guidance in semantically rich features space - h-space [18] of DMs.
+To overcome this, we propose to perform guidance in a semantically rich feature space - h-space [23] of DMs.
 Specifically, we train an Attribute Distribution Predictor (ADP) that predicts attribute distribution directly from the h-space features.
-As ADP is trained on rich and discriminative h-space features, it is - a) implemented as a linear layer, b) requires minimal training data, and c) fast training and inference.
+As ADP is trained on rich and discriminative h-space features, it - a) is implemented as a linear layer, b) requires minimal training data, and c) is fast in training and inference.
 Finally, during inference, we steer the h-space representation by matching the predictions from ADP to $\mathbf{p^{a}_{ref}}$.
@@
   1. 1.
-    A novel setting for debiasing existing DMs without retraining given a reference attribute distribution.
+    A novel setting for debiasing existing DMs without retraining, given a reference attribute distribution.
   2. 2.
@@
 Biases in Generative Models.
-While generative models like Generative Adversarial Networks and Diffusion Models have become the de-facto tools for image generation in recent times, studies show that they are not free of biases [29, 21, 34, 22, 13].
+While generative models like Generative Adversarial Networks and Diffusion Models have become the de-facto tools for image generation in recent times, studies show that they are not free of biases [35, 26, 41, 27, 15, 3].
 Perera et al.
@@
 [52] manipulate the latent space of GANs to generate balanced outputs.
+Controlling the generation with latent manipulation is easier in GANs due to its highly disentangled latent spaces in GANs [50, 34, 18].
 GANDiffFace [29], on the other hand, generates balanced synthetic for face recognition by first generating high-quality images from different demographics using GANs and then finetuning Stable Diffusion [40] using DreamBooth [42] to generate more images of such identities with different poses, expressions, etc.
@@
 Directly computing $\mathbf{p^{a}_{\theta}}$ in a closed form is intractable.
-Instead, we train an Attribute Distribution Predictor, a small network that maps the intermediate batch features of a diffusion network to an estimate of attribute distribution $\mathbf{\hat{p}^{a}_{\theta}}$.
-Figure 2: Sample guidance vs Distribution guidance.
-a) After a few steps of denoising ($t=\tau$), the generated samples have learned some discriminative features for gender.
-Sample guidance enforces a preset attribute state (male/female) to all the samples in the batch.
-This is stringent and can result in enforcing samples with dominant female features to be changed.
-However, in distribution guidance, only those samples are converted, which are easy to convert and are close to the decision boundary.
-b) Visualization of the generated samples with different guidance strengths.
-It can be observed that sample guidance is not able to change the attribute in some cases, and increasing guidance strength results in collapse.
+Instead, we train an Attribute Distribution Predictor, a linear projection that maps the intermediate batch features from h-space of a denoising network to an estimate of attribute distribution $\mathbf{\hat{p}^{a}_{\theta}}$.
 3.1 Preliminary
 Diffusion models have emerged as a powerful family of generative models trained to learn the data distribution by gradual denoising from a Gaussian distribution.
-Starting from a clean point $\mathbf{x_{0}}$, and a set of scalar values $\{\alpha_{t}\}_{t=1}^{T}$, applying $\mathbf{t}$ steps of the forward diffusion process yields a noisy data point, where $\bar{\alpha_{t}}$ = $\prod_{i=1}^{t}\alpha_{i}$
-$\mathbf{x_{t}}=\sqrt{\bar{\alpha_{t}}}\mathbf{x_{0}}+(\sqrt{1-\bar{\alpha_{t}}})\epsilon,\epsilon\approx\mathcal{N}(0,I)$ (1)
+Starting from a clean point $\mathbf{x_{0}}$, and a set of scalar values $\{\alpha_{t}\}_{t=1}^{T}$, applying $\mathbf{t}$ steps of the forward diffusion process yields a noisy data point $\mathbf{x_{t}}$, where $\bar{\alpha_{t}}$ = $\prod_{i=1}^{t}\alpha_{i}$ and
+$\mathbf{x_{t}}=\sqrt{\bar{\alpha_{t}}}\mathbf{x_{0}}+(\sqrt{1-\bar{\alpha_{t}}})\epsilon,\epsilon\approx\mathcal{N}(0,I)\vspace{-2mm}$ (1)
 A diffusion model is learned as a neural network $\epsilon_{\theta}$ that predicts the noise from given $\mathbf{x_{t}}$ and $\mathbf{t}$.
 The reverse process takes the form $q(\mathbf{x_{t-1}}|\mathbf{x_{t}},\mathbf{x_{0}})$, which is parameterized as a Gaussian distribution.
-In this work, we consider DDIM [42] sampling which first computes a clean data point $\mathbf{\hat{x}_{0}}$ and then sample $\mathbf{x_{t-1}}$ from $q(\mathbf{x_{t-1}}|\mathbf{x_{t}},\mathbf{\hat{x}_{0}})$.
+In this work, we consider DDIM [51] sampling which first computes an estimate of the clean data point $\mathbf{\hat{x}_{0}}$ and then sample $\mathbf{x_{t-1}}$ from $q(\mathbf{x_{t-1}}|\mathbf{x_{t}},\mathbf{\hat{x}_{0}})$.
 Classifier guidance is proposed to condition a diffusion model on class labels with the help of a pretrained classifier [9].
-Specifically, a classifier $f_{\phi}(c|\mathbf{x_{t}},t)$ is trained on noisy images at different noise scales, where $c$ is the class label.
-The gradients of the classifier are then used to guide the diffusion sampling process towards the class label $c$.
-Concretely, the classifier guidance performs sampling by updating the noise prediction $\epsilon_{\theta}(\mathbf{x_{t}},\mathbf{t})$.
+Specifically, a classifier $f_{\phi}(c|\mathbf{x_{t}},t)$ is trained on noisy images to predict the class label $c$.
+The gradients of the classifier are used to guide the diffusion sampling process to generate an image of the prescribed class $c$.
+Concretely, the classifier guidance performs sampling by updating the noise prediction $\epsilon_{\theta}(\mathbf{x_{t}},\mathbf{t})$ as follows:
 $\hat{\epsilon}_{\theta}(\mathbf{x_{t}},t)=\epsilon_{\theta}(\mathbf{x_{t}},t)-\sqrt{1-\alpha_{t}}\nabla_{\mathbf{x_{t}}}\log f_{\phi}(c|\mathbf{x_{t}},t)$ (2)
 3.2 Classifier guidance for debiasing
-A promising approach is to leverage pretrained attribute classifiers to guide toward balanced generation.
-Assuming a reference attribute distribution $\mathbf{p^{a}_{ref}}$ for a binary attribute $\mathbf{a}$ (e.g., gender) and corresponding attribute classifier ${f}_{\phi}$.
-We parameterize $\mathbf{p^{a}_{ref}}$ as a Bernoulli distribution with parameter $\mathbf{r}$ denoting fraction of males samples.
+A promising approach is to leverage pretrained attribute classifiers to guide towards balanced generation.
+Assuming a reference attribute distribution $\mathbf{p^{a}_{ref}}$ for a binary attribute $\mathbf{a}$ (e.g., gender) and corresponding attribute classifier ${f}_{\phi}$, we parameterize $\mathbf{p^{a}_{ref}}$ as a Bernoulli distribution with parameter $\mathbf{r}$ denoting fraction of males samples.
 To generate samples following $\mathbf{p^{a}_{ref}}$, we can randomly select $\mathbf{Nr}$ samples from the batch size of $\mathbf{N}$ and guide them towards male class using predictions from ${f}_{\phi}$.
@@
 In practice, such a sample guidance follows $\mathbf{p^{a}_{ref}}$ up to some extent, but results in inferior sample quality (see Fig.
+3b).
+Insight.1: Transforming samples close to the decision boundary is easier and results in higher quality generation.
+Remark: We performed an insightful experiment for changing gender attribute (female to male) using classifier guidance.
+We group the samples in four quantiles based on their distance from the decision boundary of a pre-trained gender classifier in Fig.
+2a).
+Next, we perform sample-based guidance over the samples from each quantile in Fig.
 2b).
-Insight. 1: Attempting to steer the generation of individual samples towards a pre-defined attribute class is overly restrictive and leads to inferior generation quality.
+The samples which are close to the decision boundary (quantiles Q1 and Q2) are easily transformed with guidance, whereas the samples away from the decision boundary (quantiles Q3 and Q4) are distorted during the guidance process.
+This is also quantified with the quantile-wise FID against a real set (of male images), where the samples from Q1/Q2 have better FID after conversion.
+Figure 2: Distribution of samples w.r.t. decision boundary
+Insight.2: Attempting to steer the generation of individual samples towards a pre-defined attribute class is overly restrictive and leads to inferior generation quality.
 Remark: Sample guidance requires enforcing a preset attribute state (male/female) to each sample of a batch during the denoising process, which is too stringent and results in distorting the outputs.
+This is particularly bad in the case when samples from quantiles Q4 are selected for transformation, as shown in Fig.
+2.
 Given an intermediate time-stamp $\tau$, the samples in the earlier stages ($t>\tau$) of reverse diffusion process are close to the noise space, resulting in the poor classifier of the corresponding stages.
 The guidance becomes effective only in the later stage of denoising ($t<\tau$).
-However, till timestep $\tau$, some facial features are already formed for the samples [25].
+However, till timestep $\tau$, some facial features are already formed [30].
 Hence, enforcing a preset attribute state at later stages ($t<\tau$) is restrictive and results in sample collapse (Fig.
-2-b)).
-For e.g., for gender attribute, if a sample has formed dominant female features till $\tau$, enforcing it to be $male$ is overly restrictive, and classifier guidance distorts the generated sample.
+3-b)).
+For e.g., for gender attribute, if a sample has formed dominant female features till $\tau$ it will be in Q4, and enforcing it to be $male$ is overly restrictive, resulting in distorted generations.
+Figure 3: Sample guidance vs Distribution guidance.
+a) After a few steps of denoising ($t=\tau$), the generated samples have learned some discriminative features for gender.
+Sample guidance randomly selects samples from the batch uniformly from all the quantiles for conversion enforcing samples with dominant female features to also change.
+However, distribution guidance majorly converts the samples close to the decision boundary (Q1/Q2), which is easy to convert.
+b) Visualization of the generated samples.
+Samples transformed with distribution guidance are natural looking without any distortion, whereas images with sample guidance suffer from distortion or unnatural appearance.
 3.3 Distribution Guidance
@@
 This can be easily integrated into the reverse diffusion process as an extension of classifier guidance by modifying Eq.2.
-Insight.2: Distribution guidance provides flexibility to batch samples to easily modify the attributes to match the defined attribute distribution.
+Insight.3: Distribution guidance provides flexibility to batch samples and transforms easier samples close to the decision boundary to match the required distribution.
 Remark: As distribution guidance does not require presetting for attribute states for each sample, it gives more flexibility during the generation as long as $\mathbf{\hat{p}^{a}_{\theta}}$ follows $\mathbf{p^{a}_{ref}}$.
-E.g., in our running example of gender attribute, only those samples will change gender after $t=\tau$, which are close to the decision boundary.
-In contrast to sample guidance, where random samples are forced to change the attribute state (Fig.
-2) resulting in inferior quality, distribution guidance is able to generate fair samples with good quality.
-Observe that distribution guidance covers the samples that are close to the decision boundary for gender classification, resulting in good outputs.
+E.g., in our running example of gender attribute, only those samples will change gender after $t=\tau$, which are close to the decision boundary (in Q1 & Q2).
+In contrast, in sample guidance, random samples are forced to change the attribute state (Fig.
+3) and are spread equally in all four quantiles, resulting in inferior quality.
+This is quantified by the fraction of samples being transformed ($f_{Qi}$) with each guidance in Fig.
+3a), where distribution guidance majorly transforms samples from Q1 & Q2 and results in fair and high-quality samples (Fig.
+3b)).
 A major design decision for implementing distribution guidance is the selection of function $g_{\psi}$.
@@
 Instead, we propose to leverage intermediate semantic features from the diffusion model itself and repurpose them for the task of attribute classification.
-This framework is super efficient as compared to image space classifiers.
+This framework is very efficient as compared to image space classifiers.
 3.4 Guidance in the H-space
@@
 The trained classifiers achieve good classification performance across multiple attributes, as shown in Fig.
-3.
+5.
 Hence, we use the trained $h-space$ classifiers to realize $g_{\psi}$.
 Specifically, we train a network Attribute Distribution Predictor that maps the batch h-space features to an estimate of attribute distribution $\mathbf{\hat{p}^{a}_{\theta}}$.
-Figure 3: Classification accuracy for linear h-space classifiers and ResNet-18 image space classifiers trained on $2K$ training examples.
-h-space classifiers are data efficient and achieve superior performance even with a linear layer
 Figure 4: Distribution guidance in the h-space.
@@
 The predicted noise is then used to update batch $x^{[1:N]}_{t-1}$ using DDIM [51].
-Insight.3: H-space guidance is extremely effective and efficient as compared to image space guidance.
+Insight.4: H-space guidance is extremely effective and efficient as compared to image space guidance.
 Remark.
@@
 In the experiments section, we compare these properties of the h-space guidance.
+Figure 5: Classification accuracy for linear h-space classifiers and ResNet-18 image space classifiers trained on $2K$ training examples.
+h-space classifiers are data efficient and achieve superior performance even with a linear layer
 4 Experiments
@@
 4.1 Evaluation Metrics
-A fair generative model is evaluated on two grounds: image quality and fairness. Therefore, we begin by discussing the metrics we use to measure these two attributes of the generated images.
+A fair generative model is evaluated on two grounds: image quality and fairness. We discuss the metrics used to measure these two aspects of the generated images below.
 Fairness.
@@
 Given an attribute $\mathbf{a}$, we assume access to a high-accuracy classifier for $\mathbf{a}$ (denoted as $\mathcal{C}_{a}$), and using the predictions from the latter, we compute the following [6, 53]:
-$|\bar{p}-\mathbb{E}_{\mathbf{x}\sim p_{\theta}(\mathbf{x})}(\mathbf{y})|_{2}$
+$||\bar{p}-\mathbb{E}_{\mathbf{x}\sim p_{\theta}(\mathbf{x})}(\mathbf{y})||_{2}$
 where $\mathbf{y}$ is the softmax output of the classifier $\mathcal{C}_{a}(\mathbf{x})$, $\bar{p}$ is a uniform vector of the same dimension as $\mathbf{y}$, $p_{\theta}$ is the distribution of the generated images.
@@
 These obtained classifiers generate high-accuracy attribute predictions, as shown in Fig.
-3.
-Further details about the dataset and classifiers are provided in a supplementary document.
+5.
+Further details about the dataset and classifiers are provided in Sec.D of supplement.
 Figure 6: Balancing generated data across gender and eyeglass attributes with various guidance strategies.
@@
 For gender, Sample Guidance is able to generate only a few males while maintaining quality.
+Latent Editing produces collapsed images for gender and fails to generate eyeglasses.
 The proposed Distribution Guidance balances (close to the ratio of $0.50:0.50$) both the attributes and generates high-quality images.
@@
 Diffusion Model Architecture.
-We evaluated our proposed debiasing method for fair generation from two state-of-the-art pretrained DMs: a) an unconditional DM, P2  [4], trained on CelebA-HQ dataset, and b) a text conditional DM StableDiffusion v1.5  [33] trained on the LAION [40] dataset.
+We evaluated our proposed debiasing method on two state-of-the-art pretrained DMs: a) an unconditional DM, P2 [5], trained on CelebA-HQ dataset, and b) a text conditional DM StableDiffusion v1.5 [40] trained on the LAION [48] dataset.
 Both these models have exceptional image generation quality; however, they have significant bias concerning the sensitive face attributes, as shown in the following subsections.
-4.3 Evaluation and Comparisons.
-We compare our proposed method against two techniques for guidance-based generations for DMs and a state-of-the-art sampling-based technique for debiasing generative models, MagNet [12].
+4.3 Baselines
+We compare our proposed method against two techniques for guidance-based generations for DMs, one Latent-based editing method [23], and a state-of-the-art sampling-based technique for debiasing generative models, MagNet [14].
 Sample guidance.
@@
 The key idea is to use the DDIM scheduler and predict the approximation of $\mathbf{\hat{x_{0}}}$ from noise image $\mathbf{x_{t}}$, and pass it through pretrained image classifier $\mathcal{C}_{A}$.
-However, this process has two shortcomings.
-Firstly, it is slow as we have to backpropagate the gradients from the image classifier.
-Experimentally, we found the h-space guidance to be $7$ times faster than the image guidance.
-Secondly, it performs poorly at the early stage due to an inaccurate approximation of $\mathbf{x_{0}}$.
-We used the same training set of $|\mathcal{D}_{clf}|=2K$ to train the h-space and image space attribute classifiers for a fair comparison.
+However, this process has two shortcomings: it is slow as it backpropagates the gradients from the image classifier (Experimentally, we found it to be $7$ times slower than h-space guidance) and it performs poorly at the early stage due to an inaccurate approximation of $\mathbf{x_{0}}$.
+We use two settings where we vary the number of images in the training set of the image space attribute classifier: (1) Using the same training set $|\mathcal{D}_{clf}|=2K$ as the h-space classifier and (2) Using the entire CelebA-HQ dataset, i.e.
+30k images.
+Latent-based Editing [23] generates images with a specific set of attributes.
+Such a technique is popularly used in debiasing GANs [37, 17] because of the well-known disentangled latent spaces of GANs that allow for such edits.
+Recent works have shown  [23] that similar semantic control is also present in the h-space of DMs and can be used for latent-based editing.
+We capitalize on this finding and perform latent editing in the h-space to generate images of desired attributes for fair generation.
+MagNet [14] is an unsupervised method enabling fair sampling from a pretrained model.
+They propose a method for uniform sampling on the image manifold to generate under-represented groups equally.
+We generated results by MagNet sampling from a StyleGAN2 [22] model trained on the FFHQ [21] dataset from their official codebase.
+Notably, as the base model for MagNet is StyleGAN2, we cannot directly compare FIDs with our DM debiased results and report random generations from StyleGAN2 as a reference.
+4.4 Main Results
+We first present the quantitative and qualitative results on debiasing single binary attributes.
+Second, we debias multiple attributes simultaneously.
+We finally present the case of debiasing attributes with multi-class labels.
 Table 1: Evaluation of balanced generation for single attribute
-Gender Race Eyeglasses Metric FD FID FD FID FD FID Random Sampling 0.178 54.59 0.334 60.01 0.251 75.21 Universal Guidance  [1] 0.193 52.10 0.377 93.42 0.189 64.55 H-Sample Guidance 0.113 51.46 0.184 56.53 0.118 57.63 H-Distribution Guidance 0.049 50.27 0.113 52.38 0.014 51.78 Random sampling* 0.307 112.28 0.463 123.97 0.276 117.83 Magnet* 0.267 91.15 0.454 97.05 0.281 106.55
+Gender Race Eyeglasses Method FD $\downarrow$ FID $\downarrow$ FD $\downarrow$ FID $\downarrow$ FD $\downarrow$ FID $\downarrow$ Random Sampling 0.178 54.59 0.334 60.01 0.251 75.21 Universal Guidance (2k) [1] 0.193 52.10 0.377 93.42 0.189 64.55 Universal Guidance (30k) [1] 0.127 48.94 0.326 58.52 0.051 78.57 Latent Editing [23] 0.001 37.40 0.214 42.69 0.330 75.04 H-Sample Guidance (ours) 0.113 51.46 0.184 56.53 0.118 57.63 H-Distribution Guidance (ours) 0.049 50.27 0.113 52.38 0.014 51.78 StyleGAN2 - Random sampling 0.307 112.28 0.463 123.97 0.276 117.83 StyleGAN2 - Magnet [14] 0.267 91.15 0.454 97.05 0.281 106.55
 Quantitative evaluation.
-We evaluate our debiasing method for the single attribute case by generating balanced distributions of individual sensitive attributes - gender, eyeglasses, and race, in Tab.
+We evaluate our debiasing method for the single attribute case by generating balanced generations of individual sensitive attributes - gender, eyeglasses, and race, in Tab.
 1.
-For the scope of this paper, we consider all these attributes as binary, and hence expect a $1:1$ ratio of the sensitive attributes in the synthesized images (for e.g.
+As these attributes are binary, the synthesized images are expected to have a $1:1$ ratio of the sensitive attributes (for e.g.
 $0.50$ fraction of males and $0.50$ fraction of females in case of gender).
@@
 4.1.
-The proposed guidance method outperforms all the methods in terms of visual quality as measured by FID and bias generation measured by FD.
+For most attributes, the proposed guidance method outperforms all the baselines in terms of visual quality as measured by FID and bias metric measured by FD.
+Although Latent editing has a better FID and FD for gender, on qualitative evaluation (as elaborated in the next section), artifacts are seen in the images (Fig.
+6).
+Moreover, this methods fails to mitigate bias in case of multiple attributes (Tab.
+3).
 Sample guidance achieves comparable FID for gender; however, higher FD indicates inferior debiasing.
-The tradeoff between FD and FID is discussed in detail in subsection 4.5.
+The tradeoff between FD and FID is discussed in Sec.A.1 of supplement.
 This supports our thesis that the distribution guidance provides enough flexibility during generation, resulting in high-quality outputs even with a high guidance scale.
+Qualitative evaluation.
+We present results for balancing gender and eyeglasses attributes in Fig.
+6.
+We randomly sample $20$ starting noise and use individual guidance methods for debiasing.
+Without guidance, the DM mostly generates female faces for the gender attribute.
+Although latent editing achieved better quantitative metrics, it produces images with artifacts and leads to collapsed results.
+Although Sample and Universal guidance increase the number of males, some images collapse.
+On the other hand, distribution guidance generates an almost equal number of males and females without affecting the generation quality.
+Moreover, all the baselines are not able to generate eyeglasses, whereas our method leads to highly balanced generations.
 Multiple attributes.
-Additionally, we apply our method for debiasing multiple attributes simultaneously in Tab.
-2.
+We apply our method for debiasing multiple attributes simultaneously in Tab.
+3.
 Specifically, given two reference distributions $\mathbf{p^{a_{1}}_{ref}}$ and $\mathbf{p^{a_{2}}_{ref}}$, we add guidance from two pretrained attribute distribution predictors $g_{\psi_{1}}$ and $g_{\psi_{2}}$.
 For this experiment, we define both the reference distribution as uniform for each attribute ($50\%$-$50\%$ splits).
-The generated results follow the reference distribution, resulting in a balanced generation across attributes.
+The generated results follow the reference, resulting in a balanced generation across attributes.
+Further analysis is provided in Sec.B of supplement.
 Table 2: Evaluation of balanced generation for multiple attribute
-Gender + Race Eyeglasses + Race Gender + Eyeglasses Metric FD $\downarrow$ FID $\downarrow$ FD $\downarrow$ FID $\downarrow$ FD $\downarrow$ FID $\downarrow$ Random Sampling 0.256 60.68 0.292 89.14 0.214 70.97 Universal Guidance 0.283 71.84 0.264 91.54 0.157 80.57 H-Sample Guidance 0.241 59.78 0.135 67.87 0.079 52.03 H-Distribution Guidance 0.075 49.91 0.101 57.46 0.057 47.45
-Comparison with MagNet [12]: MagNet is an unsupervised method enabling fair sampling from a pretrained model.
-They propose a method for uniform sampling on the image manifold to generate under-represented groups equally.
-We generated results by MagNet sampling from a StyleGAN2 [17] model trained on the FFHQ [16] dataset from their official codebase.
-We compare our method with MagNet [12] for the task of fair generation in Tab.
-1.
-We used MagNet to generate $10K$ samples and compute the FD score.
-As the base model for MagNet is StyleGAN2, we cannot compare FID directly with our DM debiased results.
-However, we can observe that MagNet achieves similar FD scores to its baseline model and is ineffective in debiasing the three sensitive attributes: gender, race, and eyeglasses.
-However, our distribution guidance significantly improves the FD metric from the baseline DM model.
-Qualitative evaluation.
-We present qualitative results for attribute balancing with respect to gender and eyeglasses in Fig.
-5.
-We randomly sample $20$ starting noise and use individual guidance methods for debiasing.
-Without guidance, the DM mostly generates female faces for the gender attribute.
-Sample guidance and universal guidance, although increasing the number of males, collapses some images.
-On the other hand, Distribution guidance generates an almost equal number of males and females without affecting the generation quality.
-Moreover, the DM hardly generates images with eyeglasses for sample and universal guidance and also without any guidance.
-Our method, on the contrary, leads to highly balanced generations.
+Gender + Race Eyeglasses + Race Gender + Eyeglasses Method FD $\downarrow$ FID $\downarrow$ FD $\downarrow$ FID $\downarrow$ FD $\downarrow$ FID $\downarrow$ Random Sampling 0.256 60.68 0.292 89.14 0.214 70.97 Latent Editing [23] 0.124 64.84 0.219 90.63 0.230 74.93 Universal Guidance (2k) [1] 0.283 71.84 0.264 91.54 0.157 80.57 H-Sample Guidance (ours) 0.241 59.78 0.135 67.87 0.079 52.03 H-Distribution Guidance (ours) 0.075 49.91 0.101 57.46 0.057 47.45
+Multi Class attributes.
+We evaluate the efficacy of our approach in balancing multi-class attributes - age and race in Tab.
+3.
+We use FFHQ [21] dataset for both attributes and obtained annotations using pretrained models as ground truth labels are unavailable.
+For age attribute, we use a pretrained VIT age classifier [39] to produce $3$ classes: Young (${<20}$ yrs), Adult (${20-60}$ yrs) and Old (${>60}$ yrs).
+For race, we use the Fairface race classifier [16] to obtain $4$ classes: White, Black, Asian, and Indian.
+Our method successfully debiases multi-class attributes and beats the random and sample guidance in all cases.
 4.5 Generating Imbalanced Distributions
-We tested our distribution guidance for generating imbalanced attribute distribution by providing skewed $\mathbf{p^{a}_{ref}}$ i) $0.20$ female and $0.80$ male ii) $0.10$ white race and $0.90$ black race.
-These two settings are extremely challenging given that male and black race are minority groups in the dataset.
+We test our distribution guidance for generating imbalanced attribute distribution by providing skewed $\mathbf{p^{a}_{ref}}$ i) $0.20$ female and $0.80$ male ii) $0.10$ white race and $0.90$ black race.
+These two settings are extremely challenging given that male and black race are minority groups in the
+Table 3: Balanced generation for multi-class attribute
+Age (3 classes) Race (4 classes) Method FD $\downarrow$ FID $\downarrow$ FD $\downarrow$ FID $\downarrow$ Random Sampling 0.256 60.68 0.292 89.14 H-Sample Guidance 0.124 64.84 0.219 90.63 H-Distribution Guidance 0.283 71.84 0.264 91.54
+Table 4: Distribution guidance for imbalanced generation
+0.20F - 0.80M 0.10W - 0.90B Method FD $\downarrow$ FID $\downarrow$ FD $\downarrow$ FID $\downarrow$ Random Sampling 0.478 72.26 0.734 77.63 H-Distribution Sampling 0.168 51.65 0.325 53.80
+dataset.
 We present qualitative results for both the settings in Fig.7, where our distribution guidance is able to generate the defined distribution with majority males.
-Note, we have binarized the race attribute as black and whites for simplicity, and hence in the generation brown race is considered under black category.
+Note, we have binarized the race attribute as black and whites for simplicity, and hence in the generation, brown race is considered under black category.
 We report quantitative metrics in Tab.
-3, where the proposed method can achieve good FD scores with the same FID.
+4, where the proposed method can achieve good FD scores with the same FID.
+Additional experimental results are tabulated in Sec.B.1 of supplement.
 Figure 7: Generating imbalanced data with two non-uniform reference distribution - i) 0.9B-0.1W, ii) 0.8M-0.2F.
@@
 Distribution guidance generates high-quality images and closely follows the skewed reference distribution.
-Table 3: Distribution guidance for imbalanced generation
-0.20F - 0.80M 0.10W - 0.90B Method FD FID FD FID Random Sampling 0.478 72.26 0.734 77.63 H-Distribution Sampling 0.168 51.65 0.325 53.80
 4.6 Ablations
-Guidance Strength $\gamma$.
-We ablate over guidance strength parameter $\gamma$ for both the distribution guidance and sample guidance in Fig.
-7.
-Increasing $\gamma$ reduces the bias (better FD) at the cost of inferior image quality (increased FID).
-Guidance strength $1500$ achieved a good tradeoff between image quality and balancing attributes.
-Further, as compared to sample guidance, the distribution guidance achieves a better tradeoff for all the guidance strengths.
-Figure 7: Ablation over guidance strength $\gamma$
+We provide ablation over the batch size here and that over the guidance scale, h-space classifier architectures and number of training examples in Sec.A of supplement.
 Batch size.
 As we approximate $\mathbf{p^{a}_{\theta}}$ with a an estimate $\mathbf{\hat{p}^{a}_{\theta}}$ over a batch of size $N$, we ablate over different values of $N$ in Tab.
-4.
+5.
 Intuitively, using a larger batch size yields a better estimate of $\mathbf{p^{a}_{\theta}}$.
-We have found that the $N=100$ works best for all our experiments.
+We have found that the $N=100$ works best for our experiments, balancing both FD and FID.
+Additionally, our model can also handle low data regime effectively.
 Table 5: Ablation over batch size for gender balancing.
-Batch size 10 25 50 75 100 FD $\downarrow$ 0.062 0.059 0.049 0.059 0.046 FID $\downarrow$ 50.54 51.51 50.86 50.98 51.64
-Figure 8: Debiasing results on stable diffusion for gender attribute.
-The proposed distribution guidance can balance the gender attributes in the neutral prompt and generalizes well to prompts involving other professions e.g.
-firefighter, doctor.
-4.6 Experiments on Stable Diffusion
+Batch size 2 4 8 10 25 50 75 100 125 150 200 FD $\downarrow$ 0.108 0.088 0.073 0.062 0.059 0.049 0.059 0.046 0.052 0.053 0.058 FID $\downarrow$ 60.12 51.86 49.80 50.54 51.51 50.86 50.98 51.64 49.91 49.81 49.74
+Figure 8: Debiasing results on stable diffusion for gender. Distribution guidance can balance the gender attributes in prompts involving other professions e.g. firefighter and doctor.
+4.7 Debiasing Text-to-Image Diffusion Models
 We implemented our distribution guidance technique for debiasing a text-to-image generation model, Stable Diffusion (SD) v1.5 [40] concerning the gender attribute in this subsection.
+We provide results on other attributes and on mitigating spurious correlation on WaterBirds [43] generation in Sec.C of supplement.
 First, we generate a dataset $\mathcal{D}$, of $10K$ images from SD, with prompts ‘a photo of a male’ and ‘a photo of a female’ to generate a labeled dataset for training h-space classifiers.
-If we use the same CelebA-HQ dataset to train the h-space classifier, the guidance is ineffective due to a significant domain shift from the SD generations.
-Next, we obtain the corresponding labeled dataset $\mathcal{D}^{h}_{clf}$ in the h-space and train a gender classifier in the h-space.
+If the h-space classifier is trained on CelebA-HQ dataset, guidance is ineffective due to a significant domain shift from the SD generations.
+Next, we obtain the corresponding labeled dataset $\mathcal{D}^{h}_{clf}$ in the h-space and train a gender classifier in the h-spaceh-space.
 We used the trained classifier for distribution guidance following Sec.
@@
 The qualitative results are shown in Fig.8, and the quantitative results are reported in Tab.
-5.
-We can observe, as compared to the baseline (random generation without guidance), our method generates fairer gender distribution with both neutral prompts and prompts mentioning professions (e.g.
-‘a photo of a firefighter’).
-Additional results on text-to-image models are present in supp.
-material.
+6.
+Baselines.
+We compare our method with the following methods.
+(1) Random sampling from SD  [40].
+(2) ITI-Gen  [62] that learns prompt embeddings for each category of the attribute given image reference sets of each category.
+It then appends these prompts during generation to produce balanced images.
+(3) Fair Diffusion  [10] that uses a lookup table to recognize the biased concept from the text input and adds scaled attribute expressions to the prompt.
+Note, these baselines are explicitly designed to debias text-conditioned diffusion model; however, our method can debias both conditional and unconditional diffusion models.
 Table 6: Balancing gender on Stable Diffusion [40] model.
-Method Gender Doctor Firefighter FD $\downarrow$ FID $\downarrow$ FD $\downarrow$ FID $\downarrow$ FD $\downarrow$ FID $\downarrow$ Random Sampling 0.317 72.37 0.355 70.11 0.235 71.86 H-Sample Guidance 0.026 70.96 0.021 68.43 0.097 70.42 H-Distribution Guidance 0.024 70.69 0.015 67.36 0.093 69.41
+Method Gender Doctor Firefighter FD $\downarrow$ FID $\downarrow$ FD $\downarrow$ FID $\downarrow$ FD $\downarrow$ FID $\downarrow$ Random Sampling 0.317 72.37 0.355 70.11 0.235 71.86 ITI-Gen  [62] 0.049 64.79 0.072 67.81 0.184 70.12 Fair Diffusion  [10] 0.227 71.22 0.035 74.37 0.036 68.33 H-Sample Guidance (ours) 0.026 70.96 0.021 68.43 0.097 70.42 H-Distribution Guidance (ours) 0.024 70.69 0.015 67.36 0.093 69.41
 4.8 Class-imbalance in attribute classification
@@
 This adds 9k images of only Black race, and the classifier trained on this balanced data performs significantly better (Tab.
-6).
+7).
 Even when gender is balanced in both the classes, we observe a significant disparity in the accuracies for Male and Female Black samples in the vanilla classifier.
 However, our proposed method helps reduce the accuracy gap between Black males and Black females.
-This shows a potential application in generating ’class’ balanced datasets to train models for other downstream tasks, which can also mitigate bias.
+This shows a potential application in generating ‘class’ balanced datasets to train models for other downstream tasks, which can also mitigate bias.
 Table 7: Group-wise accuracies of identical classifiers trained on existing and generated (balanced) data
 Black female White female Black male White male Vanilla classifier 75.76 98.96 70.29 99.93 Balanced classifier 91.04 97.33 90.72 97.68
-5 Discussion.
+5 Discussion
 Limitations.
@@
 We believe such a setting of debiasing without retraining is practical, especially in today’s era of large-scale generative models.
+Acknowledgements. This work was supported by the Kotak IISc AIML Centre (KIAC) and Meesho. Rishubh Parihar and Abhipsa Basu are supported by PMRF fellowship.
+Appendix - Balancing Act
+Appendix A Ablation Experiments
+A.1 Guidance Strength $\gamma$
+We ablate over guidance strength parameter $\gamma$ for both the distribution guidance and sample guidance in Fig.
+9.
+Increasing $\gamma$ reduces the bias (better FD) at the cost of inferior image quality (increased FID).
+Guidance strength $1500$ achieved a good tradeoff between image quality and balancing attributes.
+Further, as compared to sample guidance, the distribution guidance achieves a better tradeoff for all the guidance strengths.
+Figure 9: Ablation over guidance strength $\gamma$
+A.2 Data efficiency of h-space classifiers
+We ablate over the number of training examples used to train the h-space classifiers in Fig.
+10.
+Specifically, we train a gender-attribute classifier in image space and h-space using the same number of images.
+We use ResNet50 [11] as the image classifier and a linear head on top of the h-features as the h-space classifier.
+We pass the estimate $\hat{x}_{0}$ of the clean image at timestep $t$ to the image classifier to obtain the prediction.
+For h-space classification, we pass the ${h_{t}}$ to the linear head.
+The h-space classifiers are extremely data efficient and can achieve $>90\%$ accuracy for most of the time-steps, even when trained with just $500$ training examples.
+This shows the efficacy of guidance in the h-space, which can be done with only a few hundred training examples.
+Figure 10: Ablation over the number of training examples for classifiers.
+h-space classifiers achieve excellent performance even with only $200$ training examples.
+In contrast, ResNet50 classifiers require a large number of images to achieve similar performance.
+A.3 H-space classifier architectures
+We ablate over the different classifier architectures used for the classification in the h-space.
+We use a linear layer, an MLP and a small CNN architecture.
+The results are presented Tab.
+8.
+We report the average test accuracy over $50$ time steps, network parameters, and guidance time for generating a batch of images.
+With comparable accuracy, the linear layer has lesser parameters and performs guidance faster, and hence is used as the default classifier.
+Table 8: Ablation over the classifier architecture for h-space classification
+Classifier Avg. accuracy Guidance time $\downarrow$ # Parameters Linear 0.921 1.0x 3.2M CNN 0.908 2.5x 74M MLP 0.952 2.0x 200M
+Appendix B Multi-attribute Debiasing
+In this section, we generate balanced subgroups for a combination of the multiple attributes.
+We consider jointly balancing $2$ and $3$ sensitive attributes across all subgroups.
+Specifically, we explore the following settings: 1) Gender + Race: $0.25$ black males, $0.25$ black females, $0.25$ white males, $0.25$ white females 2) Gender + Eyeglasses: $0.25$ males with eyeglasses, $0.25$ females with eyeglasses, $0.25$ males without eyeglasses and $0.25$ females without eyeglasses.
+3) Gender + Race + Eyeglasses: $0.125$ for all the $8$ subgroups formed.
+The results are reported in Tab.
+9 where we compute the FD score with a balanced reference set and FID with the original CelebA-HQ [19] dataset.
+The FID score quantifies the visual quality, whereas the FD score accounts for the bias in the generations, as explained in Sec.4.1 (main paper).
+We do not use a reference set to compute FID as balancing across sub-groups leads to considerably less number of samples in the reference set.
+Table 9: Balancing attribute subgroups for fair generation
+Gender + Race Gender + Eyeglasses Gender + Race + Eyeglasses Method FD FID FD FID FD FID Random Generation 0.684 49.45 0.636 49.45 0.768 49.45 Sample Guidance 0.436 45.49 0.3 47.41 0.496 47.83 Distribution Guidance 0.224 45.37 0.2 45.92 0.408 43.94
+Figure 11: Debiasing results on stable diffusion for race and age attribute.
+We present random samples generated by the original stable diffusion model and with distribution guidance.
+a) Balancing across race involving different professions e.g.
+firefighter, doctor.
+b) Balancing across age attribute for neutral prompts.
+B.1 Generalization to imbalanced distributions
+Here, we present results for generating an imbalanced distribution for the subgroups by providing a skewed reference distribution.
+Such a setting is helpful in data augmentation for under-represented subgroups as shown in Sec.4.8 in the main paper.
+We take the following two settings: 1) Gender + Eyeglasses: $0.40$ males with eyeglasses, $0.10$ males without eyeglasses, $0.40$ females with eyeglasses, and $0.10$ females without eyeglasses.
+2) Gender + Race: $0.40$ black males, $0.10$ white males, $0.40$ black females and $0.10$ white females.
+These two configurations are contrary to the originally generated distribution as blacks and eyeglasses are minority groups.
+The results are provided in Tab.
+10, where we report FD with a reference set and FID score with the CelebA-HQ dataset to evaluate bias and generation quality.
+Table 10: Generating imbalance distribution across subgroups to generate more images for under-represented groups.
+Gender + Eyeglasses Gender + Race (0.40, 0.10, 0.40, 0.10) (0.40, 0.10, 0.40, 0.10) Method FD FID FD FID Random Generation 1.1 49.45 1.444 49.45 Sample Guidance 0.472 48.66 0.756 62.48 Distribution Guidance 0.38 47.68 0.464 45.51
+Appendix C Debiasing text-to-image Diffusion Model
+C.1 Social biases in facial attributes
+We extend our experiments with Stable diffusion (SD)  [40], previously outlined in Sec.
+4.6 (main paper) by implementing our distribution guidance technique for debiasing across other (race and age) attributes.
+We observe that SD generations with a neutral prompt, such as ‘a photo of a firefighter’ or ‘a photo of a doctor’, are images of people who are predominantly of white origin.
+Similarly, when prompted with ‘a photo a person’, SD majorly generates images of young people.
+We attempt to mitigate these biases with the proposed distribution guidance method following Sec.4.6.
+The qualitative results for the same are provided in Fig.
+11.
+As evident from the results, our method achieves fairness across race and age attributes using the same neutral prompts.
+We present quantitative results in Tab.
+11, where our method achieves superior FD scores computed using CLIP attribute classifier as explained in Sec.D.1.
+C.2 Background bias in bird generation
+Our method is applicable for mitigating biases that are not social as well (i.e.
+spurious correlations).
+WaterBirds [43] is a widely used synthetic dataset for demonstrating spurious correlations, which consist of images of birds across different backgrounds.
+The images of birds (landbirds and waterbirds) are collated from the CUB dataset [58], and the backgrounds from the Places dataset [63].
+The majority of the waterbirds (sea birds) are affixed with a water background, and landbirds with land backgrounds.
+We generate a dataset of $10K$ images using SD with prompts ‘photo of a land background’ and ‘photo of a water background’.
+Additional negative prompts like - ocean, water, sea, shore, river are added to generate pure land images.
+Similarly negative prompts land, mountain, sand, forest, rocks are added to generate pure water images.
+Next, we train a background h-space classifier for guidance.
+When queried with neutral prompts for landbird and waterbird, SD follows the spurious correlation in its generation.
+Specifically, when prompted for a landbird - ‘a photo of a crow’, SD predominantly generates a crow with land as the background, whereas when prompted for a waterbird - ‘a photo of a duck’, SD generates images with water in the background.
+When we apply distributional guidance, an equal number of images across both backgrounds are generated irrespective of the bird type.
+See Fig.
+12 for qualitative results.
+Figure 12: Debiasing results on stable diffusion for backgrounds while generating birds. The proposed Distribution Guidance can balance the number of birds in various backgrounds.
+Table 11: Balancing race for multiple professions and age
+Method Race-Doctor Race-Firefighter Age Random Generation 0.356 0.423 0.488 Distribution Guidance 0.191 0.186 0.194
+Appendix D Implementation details
+D.1 h-space classifiers
+Training data.
+We created a paired training of h-space features and attribute labels $\mathcal{D}^{h}_{clf}$ for training h-space attribute classifiers.
+We start with CelebA-HQ [20] dataset $\mathcal{D}_{unf}$.
+Next, we used an off-the-shelf image space attribute classifier to obtain attribute labels for $\mathcal{D}$.
+We used CLIP as a classifier for both gender and race attributes as race labels are not present in CelebA-HQ dataset.
+Specifically, we pass $\mathcal{D}_{unf}$ to the CLIP[36] image encoder and obtain its similarity with text prompts - ‘a male’ and ‘a black person’.
+The highest and lowest similarity images are then filtered to create a labeled attribute dataset $\mathcal{D}$.
+We used $|\mathcal{D}|=2000$, where $1000$ images are for the positive class, and others are for the negative class unless mentioned otherwise.
+For eyeglass attribute, we used ground truth labels from the CelebA-HQ dataset, as the predictions from CLIP were inaccurate.
+We then embed $\mathcal{D}$ into the h-space representation using DDIM[51] inversion to obtain labeled dataset $\mathcal{D}^{h}_{clf}$ to train the classifiers.
+Model architecture.
+The h-space classifiers are implemented as a single linear layer for each diffusion timestep $\mathbf{t}$.
+We used DDIM inversion with $T=49$ timesteps and obtain $\mathcal{H}^{i}=\{h^{i}_{t}\}^{t=49}_{t=0}$ as a set of $49$ h-vectors for each image $\mathbf{i}$.
+As each classifier is linear with two output neurons (positive/ negative class), they can be jointly represented as a single fully connected layer with $2T$ output neurons.
+Optimization.
+We train the h-space classifiers with the following hyper-parameters - batch size $64$, learning rate $0.001$, and for $5$ epochs on a single NVIDIA $A5000$ gpu.
+The overall training time for a single attribute classifier is 484.37s.
+D.2 Evaluation Metrics
+We created a reference set $\mathcal{D}_{ref}$ of $5K$ images following the reference attribute distribution $\mathbf{p^{a}_{ref}}$, using the attribute labels from CelebA-HQ [19].
+The ground truth labels were used for gender and eyeglasses from CelebA-HQ.
+For race, however, as the labels are not available, they were obtained from CLIP as discussed in Sec.
+D.1.
+We use $\mathcal{D}_{ref}$ to compute the reference statistics for FID.
+For FD, we compute the discrepancy of the predicted attribute distribution from the reference distribution following Sec.4.1.
+We use the Resnet-18 [11] architecture to implement the attribute classifiers trained on CelebA-HQ [19].
+This set of classifiers needs to be different from the one used to obtain $\mathcal{D}^{h}_{clf}$ for fair evaluation.